# Physics-Informed Neural Network Approach to Quark–Antiquark Color Flux Tube

### v1

@@ I Introduction
Many researchers aspire to address such problems using a forward-thinking approach—for instance, by directly solving the second-type superconducting ODE associated with the Clem ansatz, and subsequently analyzing the parameterization of the analytical solution based on lattice simulation data. While studying the solutions of such equations is inherently meaningful, the high complexity of many problems often precludes the availability of explicit analytical expressions, thus hindering forward approaches. Conversely, in QCD research, one frequently encounters inverse problems—namely, deducing physical laws from experimental observations. These problems are inherently challenging and demand sophisticated solution strategies. Various methods exist for inferring model parameters from experimental or simulated data, and recent advances in data-driven approaches have introduced increasingly modern techniques. Machine learning (ML), a modern branch of artificial intelligence (AI), is gaining traction as a powerful tool in QCD studies and applications [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] (see also the review [40] ). ML provides a data-driven framework for uncovering the behavior of complex systems. In particular, by training deep neural networks (DNNs) on empirical data, one can infer model parameters and associated physical properties. Furthermore, embedding physical knowledge—such as PDEs or ODEs and their symmetries—into DNNs can enhance the fidelity of the learning outcomes. This class of methods is known as physics-informed neural networks (PINNs) [41].
@@ II Simple model in a type II superconductor
[…]
@@ III PINN framework and PDE/IDE
We now turn to the case of handling IDEs within the PINNs framework. In this scenario, AD remains applicable for the differential operators in the equations; however, the integral operators must be incorporated through numerical techniques, specifically by discretizing the integrals (see Fig. 3). This transformation effectively recasts the IDEs into ODEs. Prior studies addressing IDEs using discretization schemes within PINNs, such as those in Refs. [60, 61], are highly instructive—particularly the auxiliary PINNs strategy proposed in Ref. [61], which reformulates complex IDEs into systems of PDEs, thereby expanding the flexibility in designing loss functions. In the present work, we adopt a discretization scheme based on Gaussian quadrature to treat Eq. (5), rewriting the integral in a discrete form as follows
[…]
@@ IV.3 Discusstions
$L_{2}\mathrm{\leavevmode\nobreak\ error}=\frac{\sqrt{\sum_{i=1}^{M}\left|E_{z,pred}^{i}-E_{z,exact}^{i}\right|^{2}}}{\sqrt{\sum_{i=1}^{M}\left|E_{z,exact}^{i}\right|^{2}}}.$ (13)
@@
We also explored different neural network configurations, such as increasing the number of hidden layers while keeping the number of neurons per layer fixed. The resulting relative $L_{2}$ errors remained nearly unchanged, indicating that the neural architecture previously adopted is effective for training the dual superconducting equation. Finally, we analyze the discrepancies in the inferred parameters. It is important to emphasize that the Clem fit employed in lattice studies is itself an analytical solution to the governing equation, and the fitted parameters accurately reproduce the color field distribution. In contrast, our approach does not rely on the analytical form of the solution; instead, the parameters are inferred solely from the equation structure and the lattice color field data through the residuals during training. Although the inferred parameters obtained via our method differ from those derived through Clem parameterization—both in the ODE and IDE scenarios—the resulting color field distributions still reasonably reproduce the discrete lattice data. As the colour field profile depends on three parameters, the analysis indicates that direct fitting is not the sole avenue for constraining their values. This is particularly evident at $d=1.19$ fm, where the fitted values of $\mu$ and $\alpha$ deviate significantly from their inferred counterparts [45], yet both parametrizations reproduce the colour field distribution in agreement with the data.
@@ Acknowledgments
This work has been supported by the National Key R&\&&D Program of China (Grant NO. 2024YFE0109800 and 2024YFE0109802).

### v2

@@ I Introduction
Many researchers aspire to address such problems using a forward-thinking approach—for instance, by directly solving the second-type superconducting ODE associated with the Clem ansatz, and subsequently analyzing the parameterization of the analytical solution based on lattice simulation data. While studying the solutions of such equations is inherently meaningful, the high complexity of many problems often precludes the availability of explicit analytical expressions, thus hindering forward approaches. Conversely, in QCD research, one frequently encounters inverse problems—namely, deducing physical laws from experimental observations. These problems are inherently challenging and demand sophisticated solution strategies. Various methods exist for inferring model parameters from experimental or simulated data, and recent advances in data-driven approaches have introduced increasingly modern techniques. Machine learning (ML), a modern branch of artificial intelligence (AI), is gaining traction as a powerful tool in QCD studies and applications [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40] (see also the review [41] ). ML provides a data-driven framework for uncovering the behavior of complex systems. In particular, by training deep neural networks (DNNs) on empirical data, one can infer model parameters and associated physical properties. Furthermore, embedding physical knowledge—such as PDEs or ODEs and their symmetries—into DNNs can enhance the fidelity of the learning outcomes. This class of methods is known as physics-informed neural networks (PINNs) [42].
@@ II Simple model in a type II superconductor
$\frac{d}{dx_{t}}E_{z}(x_{t})-\frac{f(x_{t})^{2}}{\lambda^{2}}\cdot\frac{1}{x_{t}}\int_{0}^{x_{t}}sE_{z}(s)ds=-\frac{\phi f(x_{t})^{2}}{2\pi\lambda^{2}x_{t}},$ (5)
@@ III PINN framework and PDE/IDE
We now turn to the case of handling IDEs within the PINNs framework. In this scenario, AD remains applicable for the differential operators in the equations; however, the integral operators must be incorporated through numerical techniques, specifically by discretizing the integrals (see Fig. 3)  [^2] Here,“fractional” refers to the embedding of discretized integrals into the PINN, and does not imply that our work employs fractional-order differential operators. The term“fractional” is retained only as an original label [61].. This transformation effectively recasts the IDEs into ODEs. Prior studies addressing IDEs using discretization schemes within PINNs, such as those in Refs. [61, 62], are highly instructive—particularly the auxiliary PINNs strategy proposed in Ref. [62], which reformulates complex IDEs into systems of PDEs, thereby expanding the flexibility in designing loss functions. In the present work, we adopt a discretization scheme based on Gaussian quadrature to treat Eq. (5), rewriting the integral in a discrete form as follows
$\int_{0}^{x}y(s)ds\simeq\sum_{i=1}^{n}\alpha_{i}y(s_{i}(x)),$ (8)
@@ IV.3 Discusstions
$L_{2}\mathrm{\penalty 10000\ error}=\frac{\sqrt{\sum_{i=1}^{M}\left|E_{z,pred}^{i}-E_{z,exact}^{i}\right|^{2}}}{\sqrt{\sum_{i=1}^{M}\left|E_{z,exact}^{i}\right|^{2}}}.$ (13)
@@
[…]
Finally, we analyze the discrepancies in the inferred parameters. It is important to emphasize that the Clem fit employed in lattice studies is itself an analytical solution to the governing equation, and the fitted parameters accurately reproduce the color field distribution. In contrast, our approach does not rely on the analytical form of the solution; instead, the parameters are inferred solely from the equation structure and the lattice color field data through the residuals during training. Although the inferred parameters obtained via our method differ from those derived through Clem parameterization—both in the ODE and IDE scenarios—the resulting color field distributions still reasonably reproduce the discrete lattice data. As the colour field profile depends on three parameters, the analysis indicates that direct fitting is not the sole avenue for constraining their values. This is particularly evident at $d=1.19$ fm, where the fitted values of $\mu$ and $\alpha$ deviate significantly from their inferred counterparts [46], yet both parametrizations reproduce the colour field distribution in agreement with the data.
@@ Acknowledgments
This work has been supported by the National Key R&\&D Program of China (Grant NO. 2024YFE0109800 and 2024YFE0109802).